OpenStudy (anonymous):

The velocity of a 2.4 kg particle changes from 3i-6j to -6i +4j m/s. What is its change in kinetic energy? The velocity of a 2.4 kg particle changes from 3i-6j to -6i +4j m/s. What is its change in kinetic energy? @Mathematics

OpenStudy (amistre64):

what is the formula for kinetic energy?

OpenStudy (amistre64):

1/2 mv is what im thinking but im sure im missing a ^2

OpenStudy (anonymous):

Ec = m * 1/2 v^2

OpenStudy (apoorvk):

change in kinetic energy =KE(final) - KE(initial) Ke = (1/2)mv^2

OpenStudy (anonymous):

Sorry, Ke in english LOL

OpenStudy (apoorvk):

*KE. lol

OpenStudy (amistre64):

our velocities are just the magnitudes of your vector

OpenStudy (anonymous):

oh ok so I can just take the norm of each vector an dthe difference is the answer.

OpenStudy (anonymous):

Yes, but I'm curious: could you calculate the difference of the two vectors and use that to calculate the difference in KE?